Can you help me, please

Capital Budgeting Decision

Since LX corporation is producing at full capacity, Amanda has decided to have Han examine the feasibility of a new manufacturing plant. This expansion would represent a major capital outlay for the company. A preliminary analysis of the project has been conducted at a cost of $1.6 million. This analysis determined that the new plant will require an immediate outlay of $54 million and an additional outlay of $31 million in one year. The company has received a special tax dispensation that will allow the building and equipment to be depreciated on a 20-year MACRS schedule.

Switchover to straight-line depreciation.

For Property Placed in Service after December 31, 1986

Recovery Year

3-Year (200% DB)

5-Year (200% DB)

7-Year (200% DB)

10-Year (200% DB)

15-Year (150% DB)

20-Year (150% DB)

1

33.33

20.00

14.29

10.00

5.00

3.750

2

44.45

32.00

24.49

18.00

9.50

7.219

3

14.81*

19.20

17.49

14.40

8.55

6.677

4

7.41

11.52*

12.49

11.52

7.70

6.177

5

11.52

8.93*

9.22

6.93

5.713

6

5.76

8.92

7.37

6.23

5.285

7

8.93

6.55*

5.90*

4.888

8

4.46

6.55

5.90

4.522

9

6.56

5.91

4.462*

10

6.55

5.90

4.461

11

3.28

5.91

4.462

12

5.90

4.461

13

5.91

4.462

14

5.90

4.461

15

5.91

4.462

16

2.95

4.461

17

4.462

18

4.461

19

4.462

20

4.461

21

2.231

*Switchover to straight-line depreciation.

Because of the time necessary to build the new plant, no sales will be possible for the next year. Two years from now, the company will have partial-year sales of $17 million. Sales in the following four years will be $28 million, $37 million, $40 million, and $43 million. Because the new plant will be more efficient than LX corporation's current manufacturing facilities, variable costs are expected to be 65 percent of sales, and fixed costs will be $2.4 million per year. The new plant will also require net working capital amounting to 8 percent of sales for the next year.

Han realizes that sales from the new plant will continue into the indefinite future. Because of this, he believes the cash flows after Year 5 will continue to grow at 2.5 percent indefinitely. The company's tax rate is 40 percent and the required return is 12 percent.

4) After the empirical results, Han would like to provide the recommendation to Amanda and Board of Directors, what is Han's recommendation? Amanda also wants Han to provide a sensitivity analysis and change any one of elements documented before and see what happens? For example, increase or decrease growth rate and at what level the firm can break even when NPV=0.
Answer #1

Solution:

The above question have two part answer.

Part A: Recommendations of Han for Capital Budgeting Decision.

Part B: Senstivity Analysis.

Let us answer one by one, Part A : Recommendation of Capital Budgeting Decision.

Part A:

Preliminary analysis of the project at a cost of $1.6 million is a sunk cost and will be ignored in the capital budgeting decision.

Here given the discounting factor of 12% , assuming it is a pre tax discount rate , thereby adjusting with the tax rate of 40% , the post tax rate is 12% (1-0.40)= 7.2%, we will use this post tax discount rate for discounting cash flows.

The company has received a special tax dispensation that will allow the building and equipment to be depreciated on a 20-year MACRS schedule.

Switch over to straight-line depreciation. Therefore taking the 20-Year (150% DB) for depreciation and thereby we will calculate present value of depreciation tax shield /saving. Since Tax Benefit for Depreciation is of 21 years, we will calculate the cash flows for 21 years first, then for indefinite period.

Years Sales ( In $ million) Variable Cost ( In $ Million) (65% of Sales) Contribution (In $ Million) Fixed Cost (In $ Million) Depreciation ( In $ Million) ( Refer Note) EBIT Taxes @40% EBT Depreciation ( In $ Million) ( Refer Note) Free Cash Flows Discounting Factor @ 7.2% Discounted Value (( In $ Million)
1 0 0 0 2.4 3.23 -5.63 -2.25 -7.88 3.23 -4.66 0.9328 -4.35
2 17 11.05 5.95 2.4 6.21 -2.66 -1.06 -3.72 6.21 2.49 0.8701 2.17
3 28 18.20 9.80 2.4 5.74 1.66 0.66 0.99 5.74 6.74 0.8117 5.47
4 37 24.05 12.95 2.4 5.31 5.24 2.10 3.14 5.31 8.45 0.7572 6.40
5 40 26 14 2.4 4.45 7.15 2.86 4.29 4.45 8.74 0.7063 6.17
6 43 27.95 15.05 2.4 4.55 8.10 3.24 4.86 4.55 9.41 0.6589 6.20
Since after Year 5 , here it is at the end of 6th year , the sales will be increasing at the rate of 2.5% per year , for 7th year is 43*1.025= 44.075, similarly for following years.
7 44.08 28.65 15.43 2.4 4.20 8.82 3.53 5.29 4.20 9.50 0.6146 5.84
8 45.18 29.36 15.81 2.4 3.89 9.52 3.81 5.71 3.89 9.60 0.5733 5.50
9 46.31 30.10 16.21 2.4 3.84 9.97 3.99 5.98 3.84 9.82 0.5348 5.25
10 47.46 30.85 16.61 2.4 3.84 10.38 4.15 6.23 3.84 10.06 0.4989 5.02
11 48.65 31.62 17.03 2.4 3.84 10.79 4.32 6.47 3.84 10.31 0.4654 4.80
12 49.87 32.41 17.45 2.4 3.84 11.22 4.49 6.73 3.84 10.57 0.4341 4.59
13 51.11 33.22 17.89 2.4 3.84 11.65 4.66 6.99 3.84 10.83 0.4050 4.39
14 52.39 34.05 18.34 2.4 3.84 12.10 4.84 7.26 3.84 11.10 0.3778 4.19
15 53.70 34.91 18.80 2.4 3.84 12.56 5.02 7.53 3.84 11.37 0.3524 4.01
16 55.04 35.78 19.27 2.4 3.84 13.03 5.21 7.82 3.84 11.65 0.3287 3.83
17 56.42 36.67 19.75 2.4 3.84 13.51 5.40 8.11 3.84 11.94 0.3066 3.66
18 57.83 37.59 20.24 2.4 3.84 14.00 5.60 8.40 3.84 12.24 0.2860 3.50
19 59.28 38.53 20.75 2.4 3.84 14.51 5.80 8.71 3.84 12.54 0.2668 3.35
20 60.76 39.49 21.27 2.4 3.84 15.03 6.01 9.02 3.84 12.85 0.2489 3.20
21 62.28 40.48 21.80 2.4 1.92 17.48 6.99 10.49 1.92 12.41 0.2322 2.88
Total Discounted Cash Flows 86.06

Total Discounted Cash Flows = $ 86.06 million

Determination of Present Value of Continuing Free Cash Flows

CV= Future Cash flows at year 22 (1+ growth)/Discounting factor- growth = 2.88(1.025)/0.072-0.025 = 2.952/0.047

= $62.80 million

Present Value of Continuing Free Cash flow = 62.80* Present Value Factor (7.2% , 21year ) = 62.80*0.2322= $14.58million

Total Cash Inflow = Total discounted cash flow + Present Value of Free Cash flow ( continuing) = $86.06 million + $14.58 million= $100.64 million

Present Cash outflow

Initial Cash outlay at Year 0 = $ 54 million

Add: Present Value of Additional Outlay at year 1 = $ 31*0.9328= $28.92 million

Add: Present Value of Working Capital ( as 8 % of next year sales at year 2)= 17*8%*0.8701= 1.18

Net Cash Outflow = ( 54+28.92+1.18) = $84.10 million

Net Present Value

a) Total Cash Inflow = $ 100.64 million

b) Less: Cash outflow= $ 84.10 Million

Net present Value ( a-b)= $16.54 Million

Since project NPV is positive therefore , the new manufacturing plant should be installed.

Note:

Depreciation Chart based on MACR4
Year Opening Balance (In $ Million) Addition (In $ Million) Closing Balance (In $ Million) Depreciation Rate Depreciation (In $ Million)
0 54 0 54 0 0
1 54 31 86 3.750 3.23
2 86 7.219 6.21
3 86 6.677 5.74
4 86 6.177 5.31
5 86 5.173 4.45
6 86 5.285 4.55
7 86 4.888 4.20
8 86 4.522 3.89
9 86 4.462 3.84
10 86 4.461 3.84
11 86 4.462 3.84
12 86 4.461 3.84
13 86 4.462 3.84
14 86 4.461 3.84
15 86 4.462 3.84
16 86 4.461 3.84
17 86 4.462 3.84
18 86 4.461 3.84
19 86 4.462 3.84
20 86 4.461 3.84
21 86 2.231 1.92

Part B

Sensitivity Analysis

Here NPV = $ 16.54 million

Any of the elements change whether increase or decrease, let us change the machinery value by 10 % in Year 0 and Year 1, we get increase of ($54 million * 10%+ $ 31 million *10%*0.9328 ) = $8.29 million

we get percentage fall in NPV = 8.29/16.54 *100 = 50.13%